OpenStudy (richyw):

Show that if the columns of a matrix \(Q\) form an orthonormal basis for \(C(Q)\), then \(Q^TQ=1\)

OpenStudy (richyw):

that one should be an I, the identity matrix

OpenStudy (richyw):

it's really tough to do these proofs as my textbook doesn't really do them

OpenStudy (fwizbang):

If the columns form an orthonormal basis, then Q = (e_1,e_2,e_3,....e_N) where each of the e_i is a basis vector. Since the basis is orthonormal, all the dot products obey e^T_i dot e_j = 1 if i=j, 0 otherwise. All that's left to do then is for Q^T and do the multiplications......

OpenStudy (richyw):

thanks, just throwing the brackets on your formulas! If the columns form an orthonormal basis, then \(Q = (e_1,e_2,e_3,....e_N)\) where each of the \(e_i\) is a basis vector. Since the basis is orthonormal, all the dot products obey \[e^T_i \cdot e_j = 1 \text{ if } i=j, 0 \text{ otherwise}\]. All that's left to do then is for \(Q^T\) and do the multiplications......